40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 873     Entry time: Sat Aug 23 09:39:51 2008
Author: rana, jenne 
Type: Update 
Category: PSL 
Subject: PMC Survey 
Jenne, Rana

We scoped out the PMC situation yesterday.

Summary: Not broke. UGF ~ 500 Hz. Needs some electronics work (notches, boosts, LPFs)

Ever since we swapped out the PMC because of the broken PZT of the previous one, the UGF has been
limited to a low value. This is because the notches no longer match the mechanical resonant
frequencies of the body. The old one had a resonance at 31.3 kHz which we were notching using
the LC notch on the board as well as a dangling Pomona box in the HV line to the PZT. The one
has a resonance at ~14.5 kHz which we don't yet have a notch for. Jenne has all the real numbers and
will update this entry with them.

Todo:
  • Implement the 4th order Grote low pass after the mixer.
  • Replace the AD797 with an OP27.
  • Change servo filter to have a boost (need DC gain)
  • Make a 14.5 kHz notch for the bode mode.
  • Put a 20 lb. gold-foil wrapped lead brick on the PMC.

Here's the link about the modified PMC board which we installed at LHO:
LHO PMC elog 2006
ELOG V3.1.3-